OpenStudy (butterflydreamer):

PROBABILITY question (cringe) The probability that it rains on any particular day during the 30 days of November is 0.1 Write an expression for the probability that it rains on fewer than 3 days in November.

rvc (rvc):

possible outcome ,p=0.3 since p+q=1 q=?

OpenStudy (butterflydreamer):

y = 1-p ? y = 1 - 0.3 = 0.7 ?

rvc (rvc):

correct :) Now what condition is given?

OpenStudy (butterflydreamer):

err... i haven't done probability in class so i'm kinda just guessing here LOL. condition is P ( <3 days rain) ?

OpenStudy (butterflydreamer):

i was thinking though.. wouldn't it just be something like: P (rain) = 0.1 P (no rain) = 0.9 P (<3 days rain) = P(0 days) + P (1 day) + P(2 days) ?

OpenStudy (welshfella):

yes

OpenStudy (butterflydreamer):

yeah... but i'm not sure where to go from here?

OpenStudy (welshfella):

what's the probability it will rain on 2 days you have p(0) and p(1)

OpenStudy (butterflydreamer):

hmm.